OpenStudy (rayman5766):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

OpenStudy (rayman5766):

i had put d

OpenStudy (anonymous):

what do we use to find the vairble?

OpenStudy (anonymous):

the opsite of _ is +

OpenStudy (gorv):

z+0.3=-2.4 z=-2.4-0.3=-2.7

OpenStudy (anonymous):

-0.3 + 2.4=?

OpenStudy (rayman5766):

2.1

OpenStudy (anonymous):

gorv is a mathlete so its b or a

OpenStudy (rayman5766):

ok next What is the solution to the equation? x – 1.2 = 3.6

OpenStudy (anonymous):

opss -3.0 + -2.4

OpenStudy (rayman5766):

5.4

OpenStudy (anonymous):

nvm minethe other way was right

OpenStudy (anonymous):

@mathgeek27 @mathmath333

OpenStudy (anonymous):

is positvie or negative thats the question

OpenStudy (anonymous):

c.

OpenStudy (rayman5766):

ok

OpenStudy (anonymous):

the other one i cant see